OpenStudy (anonymous):

the position of a point during the interval of time from t=0 to t=6sec is given by S=-0.5t^3+6t^2+4t. at what time does the maximum level of the point occur. what is the maximum velocity and the acceleration when the velocity is maximum

OpenStudy (anonymous):

@sweetburger @satellite73 @mathstudent55 @mathmath333 @

OpenStudy (anonymous):

\[S=-0.5t^3+6t^2+4t\] right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

at what time does the maximum level of the point occur? take the derivative, set it equal to zero and solve

OpenStudy (anonymous):

i sure hope this is a calc class, otherwise i am clueless

OpenStudy (anonymous):

yah can you help me to aarrive for the right answer

OpenStudy (anonymous):

sure the derivative is a quadratic so it can't be that hard to solve

OpenStudy (anonymous):

actually this function might be increasing on the interval from 0 to 6, largest at the endpoint did you get the deivative

OpenStudy (anonymous):

\[\frac{ dx }{ dt } = -1.5t^{2} +12 t +4\]

OpenStudy (anonymous):

is tht right?

OpenStudy (anonymous):

yeah and i think the zeros of this are outside your interval lets check i t

OpenStudy (anonymous):

how can you show to me?

OpenStudy (anonymous):

\[-1.5t^2+12t+4=0\] i would get rid of the decimal and use the quadratic formula maybe multiply by -2 and start with \[3t^2-24t-8=0\]

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

\[t=\frac{24\pm\sqrt{12^2-4\times 3\times (-8)}}{6}\]

OpenStudy (anonymous):

aka quadratic formula and a calculator too for this pretty sure the answer is 8 point something, outside your interval, so function has a max a at the right hand endpoint, namely at 6

OpenStudy (anonymous):

as a matter of fact, i am sure of it your function is increasing on the interval \([0,6]\) so it is largest at the right hand endpoint of the interval

OpenStudy (anonymous):

ok that is wrong sorry max velocity is the max of the derivative, aka the vertex

OpenStudy (anonymous):

find the vertex of \[s'(t)=-1.5t^2+12t+4\] using \(-\frac{b}{2a}\)

OpenStudy (anonymous):

to find the acceleration at that time, use the second derivative and plug in the first coordinate of the vertex of the first derivative

OpenStudy (anonymous):

let me know if i lost you anywheres there